OpenStudy (anonymous):

MEDAL! does anyone have 04.05 Solving Polynomial Equations flvs assignment? please I need it

OpenStudy (here_to_help15):

put the question :P

OpenStudy (anonymous):

ok one sec

OpenStudy (here_to_help15):

Y IS EVERY1 IN FLVS

OpenStudy (anonymous):

lol idk

OpenStudy (anonymous):

Welcome to Oakville Lake Amusement Park! As one of the new roller coaster engineers, you have been tasked with developing a roller coaster that will intertwine with existing Oakville Lake Amusement Park structures. For one of the more thrilling sections, the roller coaster will dive down in-between buildings, plummet underground, pop back up, and coast over a hill before shooting back underground. There must be three distinct points where the roller coaster crosses the x–axis. Precise measurements and attention to detail are very important. First, here is the existing map of current structures. It is important that the roller coaster does not go through the foundation of any of these structures. 1st point: ___6___ 2nd point:___-2___ 3rd point: ___-7___

OpenStudy (anonymous):

1. Using the points above as zeros, construct the polynomial function, f(x), that will be the path of your roller coaster. Show all of your work. 2.Using both fundamental Theorem and Descartes` rule of signs, prove to the construction foreman that your funtion matches your graph. Use complete sentences. 3. Solve for the y–intercept for your function, f(x), and then construct a rough graph of your roller coaster. If your y–intercept is off the graph, give the coordinates of the y–intercept.

OpenStudy (here_to_help15):

@phi u solved this problem before can u help him or her plz

OpenStudy (anonymous):

Yes please! @phi

OpenStudy (phi):

you picked your x values where f(x) is zero x= -7 x= -2 x= 6 if you add +7 to both sides for the first one, you get x+7= -7+7 or x+7= 0 do the "same thing" for the other two equations: x+2= 0 x-6= 0 now multiply all three together (x+7)(x+2)(x-6)= 0 that is your polynomial

OpenStudy (anonymous):

ok great and the number 2?

OpenStudy (anonymous):

@phi